# Transparent Networks for Multivariate Time Series

### v2

@@ Appendix E Additional visualizations
Time-step importance: Figure 9 illustrates the average importance of all time steps at the final time step. The importance exhibit a cyclical pattern of rising and falling at regular intervals, indicating that GATSM effectively captures seasonal patterns in the Rainfall dataset.
@@
Global feature contribution: Figure 10 illustrates the global behavior of features in the Rainfall dataset, with red bars indicating the density of training samples. Our findings indicate that low Max Temperature and high Min Temperature contribute to an increase in rainfall.
@@
Local time-independent feature contribution: Figure 11 shows the local time-independent feature contributions. Consistent with the global interpretation, Avg. Temperature and Min Temperature have positive correlations with rainfall, while Max Temperature has a negative correlation with rainfall.
@@
Local time-dependent feature contribution: Figure 12 shows the local time-dependent feature contributions. All features exhibit patterns similar to the local time-independent contributions. However, we found that Avg. Temperature and Min Temperature have time lags between feature values and contributions.

### v3

@@ Appendix G Visualizations
Figure 12 illustrates the average importance of all time steps at the final time step. The importance exhibit a cyclical pattern of rising and falling at regular intervals, indicating that GATSM effectively captures seasonal patterns in the Rainfall dataset.
Global feature contribution:
@@
Figure 12 illustrates the global behavior of features in the Rainfall dataset, with red bars indicating the density of training samples. Our findings indicate that low Max Temperature and high Min Temperature contribute to an increase in rainfall.
Local time-independent feature contribution:
@@
Figure 12 shows the local time-independent feature contributions. Consistent with the global interpretation, Avg. Temperature and Min Temperature have positive correlations with rainfall, while Max Temperature has a negative correlation with rainfall.
Local time-dependent feature contribution:
@@
Figure 12 shows the local time-dependent feature contributions. All features exhibit patterns similar to the local time-independent contributions. However, we found that Avg. Temperature and Min Temperature have time lags between feature values and contributions.
Mortality
Figure 13 and Figure 14 illustrates the global behavior of features in the Mortality dataset. Figure 15 and Figure 16 shows the local time-independent feature contributions to the output value. Figure 17 and Figure 18 shows the local time-dependent feature contributions to the output value. Given that the Mortality dataset consists of varying-length time series, the significance of individual time steps differs considerably for each data point. Consequently, we opted not to plot the average importance of time steps.
Figure 7: Local time-independent feature contributions.
Figure 8: Local time-dependent feature contributions.
@@
Figure 13: Global interpretations of features in the Mortality dataset (top).
Figure 14: Global interpretations of features in the Mortality dataset (bottom).
Figure 15: Local time-independent interpretations of features in the Mortality dataset (top).
Figure 16: Local time-independent interpretations of features in the Mortality dataset (bottom).
Figure 17: Local time-dependent interpretations of features in the Mortality dataset (top).
Figure 18: Local time-dependent interpretations of features in the Mortality dataset (bottom).